# Label Words are Anchors: An Information Flow Perspective for Understanding In-Context Learning

### v1

@@ 2.3.2 Results and Analysis
Figure 3a and Figure 3b show the correlation metrics of each layer for GPT2-XL and GPT-J, respectively. The result is averaged over four datasets. Firstly, the $\text{AUCROC}_{l}$ of the deep layers reaches a high score of $0.8$, indicating a strong correlation between the attention distributions on the label words of the target position and the model’s final prediction. Secondly, the cumulative contributions $R_{l}$ of the first few layers are near 0, while the score increases significantly in the middle and later layers. This phenomenon demonstrates that the classification decision of the model mainly takes place in the deep layers. From these two points, we validate that the model extracts the information from the positions of the label words to form the final prediction.
@@
3 Applications Derived from Our Anchor-Based Understanding
@@
With the insights drawn from the validated hypothesis, in this section, we propose applications to improve the effectiveness and inference efficiency. We introduce an anchor re-weighting method in § 3.1 to adaptive adjust the contribution of demonstration examples. In § 3.2, we explore a context compression technique that reduces original textual inputs to anchor hidden states for accelerating the ICL inference. Besides, using the anchor distances distribution, we also perform an analysis to better understand the errors ICL made in real-world scenarios (§ 3.3). These applications gain verify the proposed hypothesis, shedding light on new directions for future improvements of ICL.
3.1 Anchor Re-Weighting
In this part, we establish a connection between ICL and logistic regression based on the previous analysis. Inspired by such a connection, we further propose an approach for enhancing the accuracy of ICL by re-weighting the label anchors.
3.1.1 Analogy Between ICL and Logistic Regression
By approximating the ICL model as a weighted combination of classifiers and leveraging the correlation between attention distributions and final predictions, we make an analogy of ICL to ICL the logistic regression. Specifically, we demonstrate that the attention mechanisms in ICL resemble the calculation principles of logistic regression, indicating a structural resemblance between the two frameworks.
In § 2.3, we show that the output category of the model is strongly correlated with the attention values $\left(A\left(q,p_{1}\right),\ldots,A\left(q,p_{C}\right)\right)$ between the target position $q$ and the label word positions $p_{1},...,p_{C}$ at deep layers. Considering the residual mechanism of the Transformer model, the final output can be viewed as the sum of the results from previous layers. Besides, the results of each layer can be viewed as the sum of each single attention head. We can approximate the outputs of the ICL $\boldsymbol{f}$ as:
$\boldsymbol{f}\approx\sum_{l=1}^{L}\sum_{h=1}^{H}\gamma_{lh}\boldsymbol{f}_{lh},$ (6)
where $\boldsymbol{f}_{lh}$ represents the classifier approximation of the $h$th attention head in the $l$ th layer, and $\gamma_{lh}$ denotes the weight of the classifier. $\boldsymbol{f}_{lh}$ outputs a probability vector of each category for the input $x$ as follows:
$\boldsymbol{f}_{lh}(x)\approx\left(A_{l}^{h}\left(q,p_{1}\right),\ldots,A_{l}^{h}\left(q,p_{C}\right)\right).$ (7)
The approximation $\boldsymbol{f}_{lh}$ may differ from the actual one by a coefficient, but it does not affect our subsequent discussions and conclusions.
According to the calculation formula of the attention mechanism, for the $h$th head of the $l$th layer, we have:
$\text{Pr}_{\boldsymbol{f}_{lh}}(Y=i|X=x)$ (8)
$=$ $A_{l}^{h}(q,p_{i})$
$=$ $\frac{\exp(\mathbf{q}_{q}^{h}\mathbf{k}_{p_{i}}^{hT}/\sqrt{d})}{\sum_{j=1}^{N}\exp(\mathbf{q}_{q}^{h}\mathbf{k}_{j}^{hT}/\sqrt{d})},$
where $\mathbf{q}_{q}^{h}$ represents the query vector corresponding to the target position, $\mathbf{k}_{p_{i}}^{h}$ represents the key vector corresponding to the label word, and $d$ represents the dimension of the key vectors.
By defining $\mathbf{q}_{q}^{h}/\sqrt{d}=\hat{\mathbf{x}}_{lh}$ and $\mathbf{k}_{p_{i}}-\mathbf{k}_{p_{C}}=\boldsymbol{\beta}_{lh}^{i}$ (where $l$ corresponds to the layer number), we can infer that:
$\log\frac{\text{Pr}_{\boldsymbol{f}_{lh}}(Y=i|X=x)}{\text{Pr}_{\boldsymbol{f}_{lh}}(Y=C|X=x)}=\boldsymbol{\beta}_{lh}^{iT}\hat{\mathbf{x}}_{lh}.$ (9)
This is similar to the logistic regression model, where
$\log\frac{\text{Pr}_{\boldsymbol{f}}(Y=i|X=x_{t})}{\text{Pr}_{\boldsymbol{f}}(Y=C|X=x)}=\beta_{0}^{i}+\boldsymbol{\beta}^{iT}\mathbf{x}.$ (10)
$\beta_{0}^{i}$ and $\boldsymbol{\beta}^{iT}$ are learnable parameters, and $\mathbf{x}$ is the feature vector corresponding to the input.
Based on the preceding discussions, we have established an analogy between the ICL model and logistic regression, emphasizing their structural similarity. We approximate the ICL model as a combination of classifiers that employ attention distributions to generate predictions. Furthermore, we demonstrate that the computation of attention distributions exhibits similarities with the calculation principles of logistic regression. This correspondence implies that both frameworks operate on similar principles.
3.1.2 Anchor Re-Weighting Method
Inspired by the relation ICL and logistic regression, we add a learnable $\beta_{0}^{i}$ to $\boldsymbol{f}_{lh}$ in Eq. (9):
$\log\frac{\text{Pr}{\boldsymbol{f}_{lh}}(Y=i|X=x)}{\text{Pr}{\boldsymbol{f}_{lh}}(Y=C|X=x)}=\beta^{i}_{0}+\boldsymbol{\beta}_{lh}^{iT}\hat{\mathbf{x}}_{lh}.$ (11)
This is equivalent to adjust the weights of $A_{l}^{h}(q,p_{i})$, which can be expressed as
$\boldsymbol{f}_{lh}(x)\approx(\exp(\beta^{1}_{lh})A_{l}^{h}\left(q,p_{1}\right),\\ \ldots,\exp(\beta^{C}_{lh})A_{l}^{h}\left(q,p_{C}\right)),$ (12)
where $\beta^{1}_{lh},...,\beta^{C}_{lh}$ are learnable parameters.
We manipulate the attention mechanism of the model to implement this re-weighting mechanism. Please refer to the details in Appendix F.
To train a reweighting vector $\boldsymbol{\beta}=\left\{\beta^{i}_{lh}\right\}$, we use an additional training set $(\boldsymbol{X}_{train},\boldsymbol{Y}_{train})$. On the training set, we concatenate normal examples to training data to perform ICL, and optimize $\boldsymbol{\beta}$ with respect to the classification loss function $\mathcal{L}$:
$\boldsymbol{\beta}=\arg\min_{\boldsymbol{\beta}}\mathcal{L}(\boldsymbol{X}_{train},\boldsymbol{Y}_{train}).$ (13)
Metaphorically, this is equivalent to "re-weighting the anchors" in ICL. It can also be seen as an adjustment of the contribution of demonstration examples since their information has been aggregated into the anchors as suggested by our previous analysis.
3.1.3 Experiments
To verify the effect of re-weighting, we choose one sample per class as normal demonstrations, and choose $4$ extra samples per class from the task training dataset to train $\boldsymbol{\beta}$. Consistent with the setups in § 2.2, we use $5$ random seeds and report the average result. For each random seed, we fix the demonstration and sample $1000$ test samples from the test datasets. To optimize $\boldsymbol{\beta}$, we use gradient descent with the Adam optimizer Kingma and Ba (2015) with a learning rate of $0.01$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a batch size of 1 (due to memory constraint) for $10$ epochs. Due to computational resource limitations, we cannot perform training on GPT-J, so we only adopt GPT2-XL for evaluation.
We compare re-weighting with two baselines: (1) Vanilla ICL with the same demonstration (1 shot per class) (2) Vanilla ICL with the training set of $\boldsymbol{\beta}$ added to the demonstrations (5-shot per class) for a fair comparison.
3.1.4 Results
As shown in Table 2, the proposed anchor re-reweighting method significantly improves the performance of in-context learning. Particularly, the effect is remarkable on the SST-2 and EmoC datasets. Note that adding more demonstration examples for vanilla in-context learning may not bring a stable accuracy boost due to the potential noise introduced, as discussed in Zhao et al. (2021). Different from vanilla ICL which utilizes the extra examples to form a demonstration, we train a re-weighting vector $\boldsymbol{\beta}$ to adjust the contribution of different label anchors. In this way, we reduce the length of the input context and thus bring (almost) no extra cost to the inference speed. The consistent improvements of our method suggest that the re-weighting could be a better alternative to exploit the demonstration examples.
3.2 Anchor-Only Context Compression
@@
Table 2: The effect after adding parameter $\beta_{0}^{i}$. For AGNews, due to the length limit, we only include three examples for each class as a demonstration. Our Anchor Re-weighting method achieves the best performance overall tasks.
We further explore a context compression technique that reduces original textual inputs to anchor hidden states for accelerating the ICL inference.
@@ 3.2.1 Method
In § 2.3, we verify that the model’s output heavily relies on the label words, which collect information from demonstration examples during the forward propagation. We also notice that in auto-regressive language models such as GPT, the hidden state of each word depends only on the preceding words. In other words, the information aggregation process of label words is independent of the subsequent words. Therefore, we can compute the hidden states of the label words $\boldsymbol{H}=\{\{\boldsymbol{h}_{l}^{i}\}_{l=1}^{N}\}_{i=1}^{C}$ (where $h_{l}^{i}$ represents the $l$-th layer’s hidden state of the $i$-th label word in the demonstration). We concatenate $\boldsymbol{h}_{l}^{1},...,\boldsymbol{h}_{l}^{C}$ in front of the input at each layer during inference. In this way, the model can perform inference without requiring the entire input, and thus the inference can be sped up.
@@ 3.2.3 Results
Table 3: Results of different compression methods on GPT2-XL and GPT-J (averaged over SST-2, TREC, AGNews, and EmoC). Acc. denotes accuracy. The highest value excluding Vanilla ICL is marked in bold. Our method achieves the best performance among compression methods.
@@
GPT2-XL $1.1\times$ $2.5\times$ $1.5\times$ $1.4\times$
GPT-J $1.5\times$ $6\times 2.2$$2\times$ $1.9\times$
@@ 3.3 Anchor Distances for Error Diagnosis
Lastly, we perform an error analysis for ICL by utilizing the relationships of the key vectors corresponding to the label words.
@@ 3.3.1 Method
In § 2.3, we verify the correlation between the attention weights $\left(A_{l}\left(q,p_{1}\right),\ldots,A_{l}\left(q,p_{C}\right)\right)$ and the model’s output results. Here, $p_{1},...,p_{C}$ denotes the label word position indexes, and $q$ is the target position. For a single attention head, the attention score is computed as:
$A_{l}^{h}(q,p_{i})=\frac{\exp(\mathbf{q}_{q}\mathbf{k}_{p_{i}}^{T}/\sqrt{d})}{\sum_{j=1}^{N}\exp(\mathbf{q}_{q}^{h}\mathbf{k}_{j}^{T}/\sqrt{d})}.$ (14)
This implies that $A_{l}^{h}(q,p_{i})$ is influenced by $\mathbf{q}_{q}\mathbf{k}_{p_{i}}^{T}$, i.e., the similarity between the key and query vectors. Therefore, if the key vectors $\mathbf{k}$ of label words $p_{i}$ and $p_{k}$ are close, $A_{l}^{h}(q,p_{i})$ and $A_{l}^{h}(q,p_{k})$ will be relatively close for any input. Furthermore, considering the distribution of query vectors $\mathbf{q}_{q}$, we employ a PCA-like method to extract the components of the key vectors along the directions with significant variations in $\mathbf{q}_{q}$, and concatenate all heads to get feature $\hat{\mathbf{k}}$ (see Appendix H for details). The confusion between categories can then be measured by evaluating the distance between $\hat{\mathbf{k}}$:
@@
where $\text{Confusion}_{ij}^{\text{pred}}$ is a value that does not exceed 1. A larger $\text{Confusion}_{ij}^{\text{pred}}$indicates a lighter degree of confusion between label categories.
3.3.2 Settings
We select GPT2-XL and the TREC dataset, as we observe that the model exhibits significant confusion between certain categories while little confusion in others. Here we use the whole 500 samples of the TREC test set, and sample 1 demonstration per class for convenience of analysis.
@@ 3.3.3 Experiments
We measure the actual model confusion score $\text{Confusion}_{ij}$ between category $i$ and category $k$ using the AUC-ROC metric (detailed in Appendix I). The heatmap with $\text{Confusion}_{ij}^{\text{pred}}$ and $\text{Confusion}_{ij}$ is plotted for comparison.
@@ 3.3.4 Results
Figure 4: Predicted confusion matrix and real confusion matrix on TREC. We set the undefined to 11 diagonal for better visualization. Two heatmaps are similar for confusing category pairs, especially in light-color blocks.
As shown in Figure 4, the proposed approximation metric based on the anchor vectors can identify the most severe confusion case (Description-Entity) and performs reasonably well for relatively high confusion categories (Entity-Abbreviation, Description-Abbreviation). This high correlation indicates that ICL makes errors in categories with similar label anchors. Overall, this result demonstrates that our anchor-based analysis framework could become an interpretable tool for better understanding the errors of ICL.
@@ 4 Related Work
The existing literature on in-context learning analysis can be broadly divided into two streams, each focusing on different aspects. The first stream explores the influencing factors of ICL based on input perturbation, such as the order (Min et al., 2022b), the formatting (Kim et al., 2022; Wei et al., 2022), and the selection of the demonstration (Liu et al., 2022). Designing proper demonstration construction strategies (Ye et al., 2023; Li et al., 2023) and calibration techniques (Zhao et al., 2021; Min et al., 2022a) could bring clear boosts to the ICL performance. The second stream investigates the inner working mechanism of ICL through different conceptual lenses, such as making an analogy of ICL to gradient descent (von Oswald et al., 2022; Dai et al., 2022) and viewing the process of ICL as a Bayesian inference (Xie et al., 2022).

### v2

@@ 2.3.2 Results and Analysis
Figures 5a and 5b delineate correlation metrics for GPT2-XL and GPT-J, averaged across four datasets. The $\text{AUCROC}_{l}$ for deep layers approaches $0.8$, illustrating a strong correlation between the attention distributions on label words of the target position and the model’s final prediction. Moreover, shallow layers show negligible cumulative contributions ($R_{l}$), with a significant increase in middle and deep layers. These results signify the crucial role of deep layers for final prediction, validating that the model extracts information from label words in deep layers to form the final prediction.
@@
2.4 Discussion of Our Hypothesis
In § 2.2, we have affirmed that the model’s shallow layers assemble information from demonstrations via label words to form semantic representations. In § 2.3, we verify that the aforementioned aggregated information on label words is then extracted to form the final prediction in the deep layers. Recognizing the crucial function of label words in this process, we have introduced the term “Anchors” to denote them. Given the considerable role these “anchors” fulfill, we find it intuitive to design ICL improvements based on them, as elaborated in § 3.
3 Applications of Our Anchor-Based Understanding
@@
With insights from the validated hypothesis, we propose strategies to boost ICL’s accuracy and inference speed. We propose an anchor re-weighting method in § 3.1 to adjust the demonstrations’ contributions and improve accuracy. In § 3.2, we explore a context compression technique that reduces original demonstrations to anchor hidden states to speed up ICL inference. Besides, in § 3.3, we utilize anchor distances to perform an analysis to understand the errors ICL made in real-world scenarios. These approaches corroborate our hypothesis, pointing to potential paths for future ICL enhancements.
@@
Table 1: The effect after adding parameter $\beta_{0}^{i}$. For AGNews, due to the length limit, we only use three demonstrations per class. Our Anchor Re-weighting method achieves the best performance overall tasks.
3.1 Anchor Re-weighting
Based on our analysis in § 2, we draw parallels between ICL and logistic regression and propose an approach to improve ICL’s accuracy by re-weighting label anchors.
@@ 3.1.1 Method
§ 2.3 illustrates a strong correlation between the model’s output category and the attention distribution $\left(A\left(q,p_{1}\right),\ldots,A\left(q,p_{C}\right)\right)$ on label words $p_{1},...,p_{C}$ of the target position $q$ in deep layers. We can view the attention module as a classifier $\boldsymbol{f}$,
$\text{Pr}_{\boldsymbol{f}}(Y=i|X=x)$ (6)
$\approx$ $A(q,p_{i})$
$=$ $\frac{\exp(\mathbf{q}_{q}\mathbf{k}_{p_{i}}^{T}/\sqrt{d})}{\sum_{j=1}^{N}\exp(\mathbf{q}_{q}\mathbf{k}_{j}^{T}/\sqrt{d})}.$
By setting $\mathbf{q}_{q}/\sqrt{d}=\hat{\mathbf{x}}$ and $\mathbf{k}_{p_{i}}-\mathbf{k}_{p_{C}}=\boldsymbol{\beta}_{i}$, we deduce:
$\log\frac{\text{Pr}_{\boldsymbol{f}}(Y=i|X=x)}{\text{Pr}_{\boldsymbol{f}}(Y=C|X=x)}=\boldsymbol{\beta}_{i}^{T}\hat{\mathbf{x}}$. (7)
This approximates a logistic regression model where:
$\log\frac{\text{Pr}_{\boldsymbol{f}}(Y=i|X=x)}{\text{Pr}_{\boldsymbol{f}}(Y=C|X=x)}=\beta_{0}^{i}+\boldsymbol{\beta}_{i}^{T}\mathbf{x}.$ (8)
In this equation, $\beta_{0}^{i}$ and $\boldsymbol{\beta}_{i}^{T}$ are parameters that can be learned, while $\mathbf{x}$ is the input feature.
Inspired by the similarity between ICL and logistic regression, we’ve incorporated a learnable $\beta_{0}^{i}$ into Eq. (7), which is equivalent to adjusting the attention weights $A(q,p_{i})$:
$\hat{A}(q,p_{i})=\exp(\beta_{0}^{i})A(q,p_{i})$ (9)
Each $\beta_{0}^{i}$ is a learnable parameter, set uniquely for different attention heads and layers. Refer to Appendix G for more details.
To train the re-weighting vector $\boldsymbol{\beta}=\left\{\beta_{0}^{i}\right\}$, we utilize an auxiliary training set $(\boldsymbol{X}_{train},\boldsymbol{Y}_{train})$. Here, we perform ICL with normal demonstrations and optimize $\boldsymbol{\beta}$ with respect to the classification loss $\mathcal{L}$ on $(\boldsymbol{X}_{train},\boldsymbol{Y}_{train})$:
$\boldsymbol{\beta}^{\star}=\arg\min_{\boldsymbol{\beta}}\mathcal{L}(\boldsymbol{X}_{train},\boldsymbol{Y}_{train}).$ (10)
This approach can be metaphorically described as "re-weighting the anchors," leading us to term it as Anchor Re-weighting. It can also be viewed as a modification of the demonstration contributions since demonstration information has been incorporated into the anchors as suggested by our prior analysis in § 2.2. Additionally, it can be interpreted as a unique adapter variant, introducing minimal parameters while preserving most of the original model. However, it is specifically designed based on our anchor hypothesis and requires fewer parameters than traditional adapters.
3.1.2 Experiments
We choose one sample per class as normal demonstrations and choose four extra samples per class to form the auxiliary training set $(\boldsymbol{X}_{train},\boldsymbol{Y}_{train})$. The setup follows § 2.2, with results averaged over five random seeds. Owing to computational constraints, we employ GPT2-XL for evaluation, excluding GPT-J. The parameters $\left\{\beta_{0}^{i}\right\}$ are trained using gradient descent. More details can be found in Appendix H.
We compare Anchoring Re-weighting with two baselines: (1) Vanilla ICL with the same demonstration (1-shot per class) (2) Vanilla ICL, where the auxiliary training set of $\boldsymbol{\beta}$ is included as demonstrations (5-shot per class) for a fair comparison.
3.1.3 Results
As Table 1 shows, the proposed anchor re-weighting significantly enhances ICL performance, particularly on the SST-2 and EmoC datasets. Besides, adding more demonstrations for vanilla ICL may not bring a stable accuracy boost due to the potential noise introduced, as discussed in Zhao et al. (2021). Different from vanilla ICL which utilizes the extra examples to form a demonstration, we train a re-weighting vector $\boldsymbol{\beta}$ to modulate label anchor contributions. This shortens the input context and thus brings (almost) no extra cost to the inference speed. The consistent improvements of our method suggest that the re-weighting mechanism could be a better alternative to utilize demonstration examples. Furthermore, it reiterates the crucial role that anchors play in ICL.
3.2 Anchor-Only Context Compression
We further explore a context compression technique that reduces the full demonstration to anchor hidden states for accelerating ICL inference.
3.2.1 Method
In § 2.3, we find that the model output heavily relies on the label words, which collect information from the demonstrations. Given the auto-regressive nature of GPT-like models, where hidden states of tokens depend solely on preceding ones, label words’ information aggregation process is independent of subsequent words. This allows for the calculation and caching of the label word hidden states $\boldsymbol{H}=\{\{\boldsymbol{h}_{l}^{i}\}_{i=1}^{C}\}_{l=1}^{N}$ ($\boldsymbol{h}_{l}^{i}$ is the $l$-th layer’s hidden state of the $i$-th label word in the demonstration). By concatenating $\boldsymbol{h}_{l}^{1},...,\boldsymbol{h}_{l}^{C}$ at the front in each layer during inference, instead of using the full demonstration, we can speed up inference.
@@ 3.2.2 Experiments
[…]
We assess all compression methods using the label loyalty and word loyalty introduced in § 2.2, in addition to classification accuracy.
@@ 3.2.3 Results
[…]
Table 2: Results of different compression methods on GPT2-XL and GPT-J (averaged over SST-2, TREC, AGNews, and EmoC). Acc. denotes accuracy. The best results are shown in bold. Our method achieves the best compression performance.
@@
GPT2-XL $1.1\times$ $1.5\times$ $2.5\times$ $1.4\times$
GPT-J $1.5\times$ $2.2\times$ $2.9\times$ $1.9\times$
@@ 3.3 Anchor Distances for Error Diagnosis
Lastly, we perform an error analysis for ICL by examining the distances between the key vectors in the attention module that correspond to the label words.
@@ 3.3.1 Method
Our previous analysis in § 2.3 shows a strong correlation between the model output and $A(q,p_{i})$, which is determined by $\mathbf{q}_{q}\mathbf{k}_{p_{i}}^{T}$ as per Eq. 7. Should the key vectors $\mathbf{k}$ for label words $p_{i}$ and $p_{k}$ be similar, $A(q,p_{i})$ and $A(q,p_{k})$ will also likely be similar, leading to potential label confusion. Furthermore, considering the distribution of query vectors $\mathbf{q}_{q}$, we employ a PCA-like method to extract the components of the key vectors along the directions with significant variations in $\mathbf{q}_{q}$, denoted as $\hat{\mathbf{k}}$ (see Appendix J for details). We anticipate that the distances between these $\hat{\mathbf{k}}$s can correspond to the category confusion of the model, thus revealing one possible origin of ICL errors. Here, we normalize the distances to a scale of 0-1, with 0 indicating the highest degree of category confusion:
@@ 3.3.2 Experiments
We utilize the GPT2-XL model and TREC dataset, as the model displays varying confusion levels between categories on this dataset. We use all 500 samples of the TREC test set and use 1 demonstration per class for convenience of analysis.
We calculate the actual model confusion score, $\text{Confusion}_{ij}$, between category $i$ and category $k$ using the AUC-ROC metric (detailed in Appendix K). We then compare the predicted confusion score, $\text{Confusion}_{ij}^{\text{pred}}$, and the actual confusion score, $\text{Confusion}_{ij}$, via heatmaps.
@@ 3.3.3 Results
Figure 6: Predicted and real confusion matrix on TREC. We set undefined diagonals to 11 for better visualization. The heatmaps display similarity in confusing category pairs, particularly in lighter-colored blocks.
Figure 6 shows that the proposed approximation metric, $\text{Confusion}_{ij}^{\text{pred}}$, can identify the most confusing case (Description-Entity) and performs reasonably well for highly confusing categories (Entity-Abbreviation, Description-Abbreviation). This high correlation indicates that ICL makes errors in categories with similar label anchors. Overall, this result demonstrates that our anchor-based analysis framework could serve as an interpretation tool for better understanding ICL’s errors.
@@ 4 Related Work
The existing literature on in-context learning analysis can be broadly divided into two streams, each focusing on different aspects. The first stream explores the influencing factors of ICL based on input perturbation, such as the order (Min et al., 2022b), the formatting (Yoo et al., 2022; Wei et al., 2022), and the selection of the demonstration (Liu et al., 2022). Designing proper demonstration construction strategies (Ye et al., 2023; Li et al., 2023a) and calibration techniques (Zhao et al., 2021; Min et al., 2022a) could bring clear boosts to the ICL performance. The second stream investigates the inner working mechanism of ICL through different conceptual lenses, such as making an analogy of ICL to gradient descent (von Oswald et al., 2022; Dai et al., 2022) and viewing the process of ICL as a Bayesian inference (Xie et al., 2022).